Circular No. 9052

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2009 L2 (YANG-GAO) Rui Yang, Hangzhou, Zhejiang, China; and Xing Gao, Urumqi, Xin-

jiang, China, report the discovery of a new comet on several survey images (limiting mag about 15) taken by Gao in the course of the Xingming Comet Survey using a Canon 350D camera (+ 10.7-cm f/2.8 camera lens) at Mt. Nanshan — the new object identified as cometary by Yang. After posting on the Minor Planet Center’s ‘NEOCP’ webpage, numerous other CCD astrometrists have reported the object to have cometary appearance during June 16.4–16.7 UT, including E. Guido and G. Sostero (remotely using a “Global Rent-a-Scope” 25-cm reflector near Mayhill, NM, U.S.A.; ten coadded unfiltered 60-s exposures show a coma of diameter ∼ 40′′ and a tail nearly 90′′ long in p.a. 245^{o}); L. Elenin (remotely using a 20-cm refractor at Tzec Maun Observatory near Cloudcroft, NM; coma of size ∼ 36′′ with a short tail in p.a. 225^{o}); E. Prosperi and P. Camilleri (remotely using a 36cm reflector at Grove Creek Observatory, Trunkey, N.S.W., Australia; 30′′ coma with 70′′ tail in p.a. 240^{o}); M. Suzuki (Utsunomiya, Tochigi, Japan, remotely using a “Global Rent-a-Scope” 25-cm reflector at Moorook, South Australia; coma diameter ∼ 30′′); Y. Ikari (Moriyama, Shiga-ken, Japan, 26-cm reflector; coma diameter 0′.8, with a 2′.2 tail in p.a. 253^{o}); R. Ligustri (Udine, Italy; remotely using a 35-cm reflector at Moorook; fifteen co-added 60-s exposures show a coma diameter of ∼ 30′′); C. Jacques and E. Pimentel (remotely using a 25-cm reflector at Moorook; 45′′ coma and a broad tail 1′.5 long in p.a. 240^{o}).

2009 UT α_{2000} δ_{2000} Mag. Observer
June 15.80840 18^{h}16^{m}19.^{s}40 −14^{o}15′30′′.7 14.2 Yang

The available astrometry, the following preliminary elliptical orbital elements by B. G. Marsden, and an ephemeris appear on MPEC 2009-M05.

T = 2009 May 19.898 TT  ω = 345.^{o}164
e = 0.63486  Ω = 259.343 }2000.0
q = 1.28936 AU  i = 16.223
a = 3.53115 AU n^{o} = 0.148535 P = 6.64 years

COMET C/2008 Q3 (GARRADD) Visual total-magnitude estimates: May 26.39 UT, 7.6 (D. A. J. Sear-

gent, The Entrance, NSW, Australia, 10×50 binoculars; 10′ coma); June 7.90, 7.0 (A. Amorim, Florianopolis, Brazil, 20×80 binoculars; moonlight).

2009 June 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green